Sharu has 2.35 in nickels and dimes. If he has a total of thirty-two coins, how many of each coin does he have?
Answer by TimothyLamb(4379)

x = number of nickels
y = number of dimes
---
x + y = 32
5x + 10y = 235
---
put the system of linear equations into standard form
---
x + y = 32
5x + 10y = 235
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 17
y = 15
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---